OpenStudy (anonymous):

Compute the value (8) (6)

OpenStudy (anonymous):

Or (8 over 6) The answer is 28

OpenStudy (anonymous):

maybe it is \[\dbinom{8}{6}\]

OpenStudy (anonymous):

Yes idk how to input that

OpenStudy (anonymous):

\[\dbinom{8}{6}=\dbinom{8}{2}=\frac{8\times 7}{2}=4\times 7=28\]

OpenStudy (anonymous):

How does the six change to a 2?

OpenStudy (anonymous):

eight choose six is the same as eight choose two because \(6+2=8\)

OpenStudy (anonymous):

plz kindly teach what it is all about..................i don't know even a bit

OpenStudy (anonymous):

you can compute eight choose six too inf you like, but you will get the same answer, just more cancellation

OpenStudy (anonymous):

I'm having trouble understanding. How did you know how to multiply?

OpenStudy (anonymous):

\[\dbinom{8}{6}\] means the number of ways you can choose six items from a set of eight

OpenStudy (anonymous):

ok ok i got it is permutations/////////////thanks...........

OpenStudy (anonymous):

first of all you should understand that it is the same as \(\dbinom{8}{2}\) because choosing 6 people out of eight to invite to your party is the same as choosing 2 out of 8 not to invite

OpenStudy (anonymous):

But then why multiply with 7?

OpenStudy (anonymous):

so you have 8 choices for the first person, seven for the second, but order doesn't count, so you don't want to count choosing justin first and kesha second differently then kesha first and justin second. by the counting principle you get \[\dbinom{8}{2}=\frac{8\times 7}{2}\]

OpenStudy (anonymous):

eight choices for the first person, seven for the second. it is an application of the counting principle there is also a formula you can use, or you can use pascal's triangle

OpenStudy (anonymous):

the general formula is \[\dbinom{n}{k}=\frac{n!}{k!(n-k)!}\] but you don't really want to ever use it

OpenStudy (anonymous):

for example you could write \[\dbinom{8}{6}=\frac{8!}{6!(8-6)!}=\frac{8!}{6!2!}=\frac{8\times 7}{2}\]

OpenStudy (anonymous):

I'm sorry I'm having trouble with how the 7 came into the equation

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

i have say 8 letters and i want to know how many ways i can choose any two of them lets call the letters \(\{a,b,c,d,e,f,g,h\}\)

OpenStudy (anonymous):

so have two slots to fill ( __ , __ ) in which to put the two chosen letters. there are 8 choices for the first letter. now that i have chosen one of the eight letters, there are only seven choices left for the second letter, because one has already been chosen

OpenStudy (anonymous):

by the counting principle, if there are 8 ways to do one thing, and 7 ways to do another, there are \(8\times 7\) ways to do them together

OpenStudy (anonymous):

but i do not want to count (a, b) and (b, a) as two separate choices, because the order i choose them in makes no difference. in other words i can't tell the difference between (a,b) and (b,a) so i have counted everything twice therefore, i have to divide by 2 to get the answer i need, i.e. \[\dbinom{8}{2}=\frac{8\times 7}{2}\]

OpenStudy (anonymous):

is it clear where the 7 came from? eight choices for the first letter, then 7 for the second because one is already chosen

OpenStudy (anonymous):

Yes I understand now!!!! But one last thing, how is 6 thee same as 2?

OpenStudy (anonymous):

Is it from the n-k?

OpenStudy (anonymous):

lets say you have six eggs to throw at cars in a parking lot, but there are 8 cars in the lot all together you can either decide which six out of the eight cars to throw eggs at (the number of ways to do this is \(\binom{8}{6}\) or you can choose two of the cars not to throw eggs at. the number of ways to do that is \(\binom{8}{2}\)

OpenStudy (anonymous):

it is the same problem exactly this shows for example that \(\dbinom{12}{4}=\dbinom{12}{8}\)

OpenStudy (anonymous):

and yes, if you want to appeal to the formula, it is from the \(k!(n-k)!\) in the denominator

OpenStudy (anonymous):

for example \[\dbinom{12}{4}=\frac{12!}{4!(12-4)!}=\frac{12!}{4!8!}=\frac{12!}{(12-8)!8!}=\dbinom{12}{8}\]

OpenStudy (anonymous):

in math you can say \[\dbinom{n}{k}=\dbinom{n}{n-k}\]

OpenStudy (anonymous):

moral of the story is pick the one that is easiest

OpenStudy (anonymous):

I understand it better! Thank you for taking the time to help me!

OpenStudy (anonymous):

yw